Calculate the concatenated transformation matrix for the following operations performed in the sequence as below:

  1. Translation by 4 and 5 units along X and Y axis
  2. Change of scale by 2 units in X direction and 4 units in Y direction

iii Rotation by 60° in CCW direction about Z axis passing through the point (4, 4). Find new coordinates when the transformation is carried out on a triangle ABC with A (4, 4), B (8, 4) and C (6, 8).
Answer #1

consider ,y) a point in co-ordinate system After Step I point becomer x-4 & 4.5 After Step 2 point becomes - 4 & 4-5 z 4 0=6014 - а + 2 x, - - - я, 9, 8, 4 - х, 9 - 1 - ) у 6, 8 x, y = 41 343 , 2 1